OpenStudy (anonymous):

evaluate the improper integral: ∫e^(-x) dx (from 0 to ∞)

OpenStudy (zarkon):

1

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

i kept getting -1

OpenStudy (zarkon):

\[\lim_{t\to\infty}\left.\int\limits_{0}^{t}e^{-x}dx=\lim_{t\to\infty}-e^{-x}\right|_{0}^{t}\] \[=\lim_{t\to\infty}-e^{-t}+e^0=0+1=1\]

OpenStudy (zarkon):

ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait why is -e^-t = 1?

OpenStudy (zarkon):

-e^-t is not equal to 1

OpenStudy (anonymous):

i know it's like -e/t...

OpenStudy (anonymous):

u said 1 + 0 = 1

OpenStudy (zarkon):

\[\lim_{t\to\infty}-e^{-t}=\lim_{t\to\infty}-\frac{1}{e^{t}}=0\]

OpenStudy (anonymous):

thanks!